OpenStudy (anonymous):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (campbell_st):

well I'll do part 1 and you can do the rest point slope form is \[y - y_{1} = m (x - x_{1})\] so substituting you have y - (-2) = 2(x - 3) which becomes y + 2 = 2(x -3) hope this helps. for the slope intercept form you want an equation in the form y = mx + b just distribute the right hand side and then subtract 2